A ball of mass 1g and charge 10(—8) C moves from a point A where potential is 600V to a point B (potential is zero). Velocity of the ball at point B is 20cm/s. Velocity of the ball at point A is:
 (plz answer with required steps)

Grade:

2 Answers

Arun
5 years ago
Dear student
 
As we know
 
Electrostatic field force is conservative.
 
600 * 10^-8 + 1/2 mv² = 0  + 1/2 m * 0.2²
 
Hence Va = 0.167 m/sec
 
 
Regards
Arun
Kushagra Madhukar
3 years ago
Hello student

As we know only the electrostatic force is acting on the charged body, which is a conservative force.
Hence, the principle of conservation of Mechanical Energy holds true.

Ki + Ui = Kf + Uf
½ mvA2 + qVA = ½ mvB2 + qVB
½ x (0.001) x vA2 + 10-8 x 600 = ½ x (0.001) x (0.2)2 + 0

On solving, we get,
vA = 0.1673 m/s = 16.73 cm/s

Hope it helps
Regards,
Kushagra